OpenStudy (anonymous):

R(x)=x^2 + 14x + 24/ x^2 + 13x + 12 find the horizontal or oblique asymptotes

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

y = 1 right?

OpenStudy (anonymous):

no oblique asymptotes since the degree of the numerator is the same as the degree of the denominator. horizontal is y = 1, the ratio of the leading coefficients

OpenStudy (anonymous):

vertical asymptotes set denominator = 0 and solve

OpenStudy (anonymous):

\[x^2+13x+12=0\] \[(x+12)(x+1)=0\] \[x=-12\] \[x=1\] are your vertical asymptotes

OpenStudy (anonymous):

thinking of course of \[x=-12\] as the vertical line , not the number

OpenStudy (anonymous):

and of course i made a typo. should be \[x=-1\] for the other one

OpenStudy (anonymous):

amistre, i got the bike!

OpenStudy (anonymous):

how do you show the steps for the above problem

OpenStudy (amistre64):

do those zeros have any tops that cancel?

OpenStudy (amistre64):

yay!! ... i got that pitman arm that needs replaced....

OpenStudy (anonymous):

good question. if so i am dead wrong \[\frac{x^2+14x+24}{x^2+13x+12}\] \[=\frac{(x+12)(x+2)}{(x+12)(x+1)}=\frac{x+2}{x+1}\]

OpenStudy (anonymous):

so i was completely wrong. horizontal asymptote still y = 1 vertical only at x = -1 and a hole at x = -12 sorry

OpenStudy (anonymous):

thanks amistre for setting me straight!

OpenStudy (anonymous):

question is what kind of moron would write \[\frac{x^2+14x+24}{x^2+13x+12}\] when the just meant \[\frac{x+2}{x+1}\]?

OpenStudy (amistre64):

lol .... sometimes that data follows the rational expression f(x)/g(x); but the graph of it can be simplified by (x+2)/(x+1)

OpenStudy (amistre64):

of course with the hole at -12